OpenStudy (idealist10):

Use Euler's method with step sizes h=0.1, h=0.05, and h=0.025 to find approximate values of the solution of the initial-value problem y'+3y=7e^(4x), y(0)=2 at x=0, 0.1, 0.2, 0.3, ..., 1.0.

OpenStudy (anonymous):

\[y'=f(x,y)=7e^{4x}-3y\] Fill in the table using \(y_{n+1}=y_n+hf(x_n,y_n)\) with \(x_{n+1}=x_n+nh\) and \(h=0.1\) for \(0\le n\le10\). \[\begin{array}{c|c|c|c} n&x_n&y_n&f(x_n,y_n)\\ \hline 0&0&2&7e^0-3(2)=1\\ 1&0+0.1=0.1&2+0.1(1)=2.1&7e^{0.4}-3(2.1)\approx 4.14277\\ 2&0+2(0.1)=0.2&2.1+0.1(4.14277)\approx 2.51428&\cdots\\ \vdots&\vdots&\vdots&\vdots\\ 10&0+10(0.1)=1&\cdots&\cdots \end{array}\] If you follow this format for \(h=0.05\) and \(h=0.025\), you'll obviously have to take more steps. Tedious? Yes.

OpenStudy (idealist10):

@SithsAndGiggles So what are the answers?

OpenStudy (anonymous):

At \(x=0\), the approximate solution is the initial condition, \(y=0\). At \(x=0.1\), i.e. the first step using \(h=0.1\), gives an approximation of \(y=2.1\). For this particular step size, each successive \(y_n\) will be the approximate solutions the question wants.

OpenStudy (anonymous):

You can try writing a script/program to do the recurrence for you, but that takes some technical coding knowledge. There are a few of these available freely on the web. This one seems to work pretty well, and agrees with (at the very least) all the points I worked out. http://www.mathscoop.com/calculus/differential-equations/euler-method-calculator.php